# Revisiting the Power of Prompt for Visual Tuning

### v2

@@ 5 Discussion and Conclusion
Conclusion. Our work proposes a simple and novel method (termed SPT) to improve visual prompt tuning. Specifically, we observe that the mutual information between prompt tokens and image patch tokens gradually increases during the prompt tuning process of VPT (blue curves in Fig. 2), and propose the idea of constructing prompts based on inferred token prototypes. This idea can make the prompt tokens and patch tokens share larger mutual information at the beginning of training (red curves in Fig. 2), which benefits the rapid convergence of the model and achieve higher accuracy.
Our method delivers substantial performance improvements, especially for self-supervised pre-trained models. For instance, after MAE pre-training (Table. 2), our method improves accuracy by up to 10$\sim$30compared to VPT, and outperforms Full fine-tuning 19 out of 24 cases while using less than 0.4% of learnable parameters. Significant performance improvements are also achieved for MoCo-V3 and supervised pre-training.
We conduct sufficient ablation experiments and analyses. The results show that our method, SPT, exhibits better than VPT (a strong baseline), e.g., SPT is robust to prompt length and achieves slight gains with increasing prompt length (Fig. 3c); Also, SPT scales well with model capacity (Fig. 3d). These characteristics still work well with supervised pre-trained backbone (Appendix Fig. S1).
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments and suggestions. This work was supported by the National Key R&D Program of China under Grant No. (2022YFF0608000, 2022ZD0119004), in part by the National Natural Science Foundation of China under Grant No. (62106235, 62376206, U21B2048, 62322605, 62293543).
Impact Statements

### v3

@@ 5 Discussion and Conclusion
Conclusion. Our work proposes a simple and novel method (termed SPT) to improve visual prompt tuning. Specifically, we observe that the mutual information between prompt tokens and image patch tokens gradually increases during the prompt tuning process of VPT, and propose the idea of constructing prompts based on inferred token prototypes. This idea can make the prompt tokens and patch tokens share larger mutual information at the beginning of training, which benefits the rapid convergence of the model and achieve higher accuracy. Our method delivers substantial performance improvements, especially for self-supervised pre-trained models. For instance, after MAE pre-training, our method improves accuracy by up to 10$\sim$30compared to VPT, and outperforms Full fine-tuning 19 out of 24 cases while using less than 0.4% of learnable parameters. Significant performance improvements are also achieved for MoCo-V3 and supervised pre-training. We conduct sufficient ablation experiments and analyses. The results show that our method, SPT, exhibits better than VPT (a strong baseline), e.g., SPT is robust to prompt length and achieves slight gains with increasing prompt length; Also, SPT scales well with model capacity. These characteristics still work well with supervised pre-trained backbone.
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments and suggestions. This work was supported by the National Key R&D Program of China under Grant No. (2022YFF0608000, 2022ZD0119004), by the Fundamental Research Funds for the Central Universities, in part by the National Natural Science Foundation of China under Grant No. (72188101, 62106235, 62376206, U21B2048, 62322605, 62293543).
Impact Statement